Investment of $9,000. Part at 3%, part at 4% (simple interest over 1 year). How much was invested in each account if the interest earned in the 3% account was $32 greater than the interest earned in the 4% account?
PRT=I

Answer by richwmiller(17219)

x+y=9000,
0.03*x-0.04*y=32
x=9000-y
0.03*(9000-y)-0.04*y=32
270-0.03y-0.04*y=32
-0.07*y=-238
y=3400 at 4%
x=9000-y
x=5600 at 3%
check
0.03*5600-0.04*3400=32
168-136=32
32=32
ok